OpenStudy (anonymous):

Determine the standard form of the equation of the line that passes through (7, -1) and (3, 9). a. 5x +2y= 33 b. 2x+5y = 33 c. 5x-2y=33 d. 5x+ 2y=-33

hartnn (hartnn):

can u find slope first ?

OpenStudy (anonymous):

umm hold on yeah

OpenStudy (anonymous):

-5/2=-2.5? That's what I got but I get stuck after getting the slope

hartnn (hartnn):

the equation of line can be written as y-y1=m(x-x1) since u have slope, u just need one point, (x1,y1) u have 2 points, take any 1 of those as (x1,y1) and substitute in y-y1=m(x-x1)

OpenStudy (anonymous):

So the answer is d?

hartnn (hartnn):

did u guess that ?

OpenStudy (anonymous):

No that's the answer I got from the beginning but I was unsure....

hartnn (hartnn):

yup, its D

OpenStudy (anonymous):

Well that's good to know that I got the right answer!!!!! Thanks!

hartnn (hartnn):

welcome :)